OpenStudy (anonymous):

How many combinations are there of the five letters j, k, l, m, and n using four letters at a time

OpenStudy (anonymous):

\[nCr = \frac{n!}{(n - r)!r!}\]\[5C4 = \frac{5!}{(5 - 4)!4!}\]

OpenStudy (anonymous):

\[5C4 = \space?\]

OpenStudy (anonymous):

whattttt

OpenStudy (anonymous):

Have you learned about factorial?